OpenStudy (kkutie7):

This is part of a "short" problem set that I just don't understand. My work below. 1.70 x 10-5 M ML was found to generate a detector response of 426 mV at 532 nm. When a non-absorbing species was placed into the cuvette, the detector response measured was 765 mV at 532 nm. Calculate molar absorption.

OpenStudy (kkutie7):

My idea: \[A=\epsilon bC=-log\frac{P}{P_{0}}\] \[-log\frac{426 mV}{765 mV}=\epsilon (1cm)(1.70*10^{-5}M)\] \[\epsilon =\frac{-log\frac{426mV}{765mV}}{(1cm)(1.70*10^{-5}M)}\]

OpenStudy (kkutie7):

Hey @dan815 am I thinking about this right?

OpenStudy (kkutie7):

@imqwerty

OpenStudy (kkutie7):

@Kainui

OpenStudy (kainui):

Yeah this is correct, I was just checking cause it's been a while. Since power is related to current and voltage by P=IE the fact that you were given the voltage (E) in milivolts is the same as if you were given the power (P) since the current is a constant throughout the experiment I believe. Specifically this substitution is safe: (if this is your concern) \[\log \frac{P}{P_0} = \log \frac{IE}{IE_0} = \log \frac{E}{E_0}\]

OpenStudy (kkutie7):

I got a very large number and I was worried about it \[14955.99M^{-1}cm^{-1}\]

OpenStudy (kkutie7):

After this I have to find this: If the uncomplexed metal (Mn+) also absorbs weakly at 532 nm (with  = 575 M-1 cm-1 ) what is the concentration of ML in a solution containing ML + 0.0916 mM Mn+ if the detector response for the solution is 169 mV?

OpenStudy (kainui):

Hmm I haven't done much with molar absorptivity in a long while so I can't be very helpful I gotta go review. I think you're right though, that value did seem pretty high since it did have a pretty high transmittance, it doesn't make sense that it would have high absorptivity.

OpenStudy (kkutie7):

All I know is that it has to be complicated because my Professor doesn't like simple and I'm confused

OpenStudy (kainui):

Gimme a little time, I should probably know this cause I have chemistry degree so I'm gonna read some stuff but I might get distracted on side tangents and take an hour or so haha

OpenStudy (kkutie7):

As Long as you can help be by tomorrow I'll be very grateful

OpenStudy (kkutie7):

@TheProfessor35 please help me

OpenStudy (kkutie7):

@Whitemonsterbunny17 do you know how to do this?